OpenStudy (anonymous):

Which expression is the simplified form of sqrt r^49

OpenStudy (anonymous):

7

OpenStudy (anonymous):

is it \[r^7 \sqrt{r}\]

jimthompson5910 (jim_thompson5910):

In general \[\Large \sqrt{x^k} = x^{\frac{k}{2}}\] if k is even OR \[\Large \sqrt{x^k} = x^{\frac{k-1}{2}}\sqrt{x}\] if k is odd

OpenStudy (anonymous):

wait that confused me a lil

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

the k/2 part

jimthompson5910 (jim_thompson5910):

Here are 2 examples: Even Example: If k = 8, then \[\Large \sqrt{x^k} = x^{\frac{k}{2}}\] \[\Large \sqrt{x^8} = x^{\frac{8}{2}}\] \[\Large \sqrt{x^8} = x^{4}\] Odd Example: If k = 13, then \[\Large \sqrt{x^k} = x^{\frac{k-1}{2}}\sqrt{x}\] \[\Large \sqrt{x^{13}} = x^{\frac{13-1}{2}}\sqrt{x}\] \[\Large \sqrt{x^{13}} = x^{\frac{12}{2}}\sqrt{x}\] \[\Large \sqrt{x^{13}} = x^{6}\sqrt{x}\]

jimthompson5910 (jim_thompson5910):

The k/2 means take whatever number k is and cut it in half

jimthompson5910 (jim_thompson5910):

Like the example above shows

OpenStudy (anonymous):

o ok i see